OpenStudy (smurfy14):

x+√(31-9x)=5 solve the equation

OpenStudy (anonymous):

31-9x=(5-x)^2=25-10x+x^2 x^2-x-6=0 (x-3)(x+2)=0 x=3 or x=-2

OpenStudy (smurfy14):

ok thanks! i made a mistake when i was working it :P